OpenStudy (anonymous):

graph each function y= | x - 4 | y= | x | + 3

OpenStudy (anonymous):

wait ...

OpenStudy (anonymous):

 first

OpenStudy (anonymous):

it is shifted 4 units towards right to the graph of y=mod(x) and upward only

OpenStudy (anonymous):

 second

OpenStudy (anonymous):

upwad here means touching x axis and above it

OpenStudy (anonymous):

for second graph...................y=mod(x)+3 the graph is shifted upward by 3 units in comparison to y=mod(x)

OpenStudy (anonymous):

ur welcome....

OpenStudy (anonymous):

no its necessary..................they are just for more clarity

OpenStudy (anonymous):

sorry i was to write no its not necessary

OpenStudy (anonymous):

you know there are 24 transformations possible for y=mod(x) curve

OpenStudy (anonymous):

ya............its shifted 3 units upward

OpenStudy (anonymous):

right,...

OpenStudy (anonymous):

any other query?

OpenStudy (anonymous):

its y=mx+c where m represents the slope and c is the y- intercept ant (x,y )the point

OpenStudy (anonymous):

no pro.